OpenStudy (anonymous):

Justin is looking at two accounts in which to invest his money. If he invests $25,000 in the first account and $5000 in the second account, then his total investment of $30,000 will receive $1700 in interest at the end of one year. If he invests $27,000 in the first account and 3000 in the second account, then his total investment of $30,000 will receive $1620 in interest at the end of one year. For each account, what percentage of the investment is received in interest after one year?

hero (hero):

Will post a solution soon.

hero (hero):

1. Set Up Variables x = 1st account percent interest y = 2nd account percent interest 2. Setup Equations 1st Possibility Total interest received = 25,000x + 5,000y = 1700 2nd Possibility: Total interest received = 27,000x + 3000y = 1620 3. Solve System Of Equations 25,000x + 5000y = 1700 27,000x + 3000y = 1620 (a) Isolate y in both equations: y = (1700 - 25,000x)/5000 y = (1620 - 27,000x)/3000 (b) Set y = y and cross-multiply: 3000(1700 - 25,000x) = 5000(1620 - 27,000x) (c) Reduce both fractions by factor of 1/100: 30(17 - 250x) = 50(16.20 - 270x) (d) Continue solving for x: 510 - 7500x = 810 - 13500x 13500x - 7500x = 810 - 510 6000x = 300 x = 300/6000 x = 3/60 x = 1/20 x = .05 4. Substitute x back into one of the original equations to find y: y = .09 5. Solution: Therefore, for each account, after one year, the percent investment received will be 5% and 9% respectively.